RFBG Commissioning with Adaptive Secondary - 20 March 2010 UT

Night 1: Testing TCS Build AO3

Observer: JHill, JStorm (LBTO)
Telescope Operator: WWack
Software: NCushing, TLeibold, TEdgin (Tucson)
Instrument Support::
Telescope Support: JUrban
Adaptive Support:: LFini, MXompero (LBTO)


We completed the closed-dome testing of TCS software Build AO3. A few minor issues were fixed along the way. Paul has a bug in OSS related to waiting for DX M3 to be in position to fix on Monday.

The software test report is at Software.SWTestReports.

SX M3 seems to have a short in one of its cables.

Sky is foggy and we are waiting for it to clear. Finally opened at 08:54 UT

In the meantime there are serious problems with the shell. Marco et al are working on it. They should get telescope priority during the day if the need it. (Marco will need to sleep at some point.)

We started with strong coma (of biblical proportion). We adjusted it with tip-tilt of M2. But we had to make tip and tilt changes of order 600 arcsec. See Jesper's photo.

Despite a peculiar shape on the shell. We managed to point at a couple stars and guide on them. This means that many things on the RFBG side are working.

Something is stuck between M2 hexapod and PSFR starting at 12:12 that we can't figure out. It worked fine earlier in the night.


Closed-Dome Tests of Build AO3 (joined in progress)

Working as IRTC@LFBG

01:00 Problems related to sending presets from IRAF. The SetStars command does not always get the new target. Sometimes, even the preset is succeeding, but with the telescope still at the previous target.

01:49 DJT sends pointto, but telescope didn't get the new target.

01:50 Repeating that several times. It turned out to be a peculiar bug in the pointto.cl script. DThompson fixed it.

Pause for dinner, LN2, etc. Sky outside is clearing, but with light snow flurries.

03:11 pointto StoneEagw gs=0 TelM="GUIDE" (failed in PARALLACTIC mode)

03:12 pointto StoneEagw gs=0 TelM="GUIDE" (succeeds in POSITION mode)s

03:14 pointto StoneEagw gs=15 TelM="GUIDE"

03:18 takepic 1

03:19 takeseq 0.1 20

03:25 Changing Authorization to Test M3 motions, but SX M3 still needs to be connected.

03:35 10 minutes to plug the three SX M3 cables (Box1) back into the M3 UMAC.

Sat Mar 20 03:33:15.565 2010 TCS authorize alert for right IRTC at bentGregorianFront focus: [2] - Command timeout - -13107: Arbitrator interface generic error

See Issue 2614. There seems to be a short in the Left M3 cables. The Left cables are disconnected, and we are using M3 UMAC on right side only.

Working as IRTC@RFBG

PCS is toggled to Right.

04:54 Torsten sends a Right side preset. But Paul is initializing M3 or something with OSS.

05:04 Authorize IRTC @ RFBG

05:07 Torsten sends TRACK preset on right side. Luca and Marco are checking for proper behaviour which started a long discussion about when a command to flatten the secondary should be needed.

05:18 Authorize as LBTI @ RFBG - OSS moves DX M3, but fails to recognize that it is done moving.

05:42 pointto StoneEagw TRACK gs=0 - hangs waiting for collimation of M3. This is Issue 2615.

Sat Mar 20 05:48:47.850 2010 Preset alert for right IRTC mode TRACK: dx.terc: [adjustCurrFocalStation.moveMirrorAbs.waitForInPositionAll] Wait for ''in position'' timed out

Change M3 to simulation mode. M3 was left at -1087.3 +18 0 -43.18 Restart OSS.

05:54 pointto StoneIagw TRACK gs=0

05:58 pointto BS9152 TRACK gs=0

06:03 pointto BS9145 TRACK gs=0

06:04 pointto StoneEagw TRACK gs=0 PA=0

Luca says "Because of a problem in the state machine, we are unable to make a repeatable sequence of good (safe) failures."

06:xx pointto StoneEagw TRACK gs=0 PA=120 (but John got the sign wrong)

06:24 pointto StoneEagw gs=0 TelMode='TRACK' PA=240 long way - successfully went from rotator angle 127 to 367.

Reconfigure Telescope Swing Arms for Bent Front Gregorian observing.

06:40 Move to zenith to reconfigure to LFBG. And since DAshby was reading his email, we'll reconfigure to RFBG too. We needed Dave to adjust the balance trim when we changed the right side.

Finally after 24 years, I'm able to deploy two secondary mirrors simultaneously!!!

07:08 Reconfiguration complete. Two sides only take a few minutes longer than one side.

07:48 Authorize IRTC on both sides, and find a couple errors (Apparently the testing is leaky.) See Issue 2618.

08:02 Open vent doors to cool down the chamber. Both the outside temperature and the dewpoint have dropped 5 degC in the past 90 minutes. Sky seems clear, but we are still closed for humidity at 94%.

08:20 Humidity has dropped, and sky is clear, so we are ready to open. But Marco is having a problem with the adaptive secondary and a group of actuators at the outer edge. These misbehaving actuators are preventing him from applying the proper flat shape.

register IRTC on Left side. Verify that PCS is set to left. But we failed to restart PCS, and we aren't sure who pressed the left button.

08:45 go to zenith to reset the shell


08:54 Open with telescope at EL=30. T=-8.0 D=-14.1 Wind moderate but not well known. Sky clear.

This is still Build AO3 verification on the SX side

08:58 pointto M5_1332 gs=0 TelM="ACQUIRE" PAM- - some issue with LFBG rotator. Preset times out.

09:05 pointto M5_1332 gs=0 TelM="ACQUIRE" PAM- - all the polys are going to the right rotator and not the left one.

09:06 stop and restart PCS. Cancel the hung preset.

09:09 pointto M5_1332 gs=0 TelM="ACQUIRE" PAM- - found a star. Adjust IE=-69 CA=-13 at EL=80.

09:15 pointto M5_0519 gs=0 TelM="ACQUIRE" PAM- - found a star Adjust IA=-280

09:19 pointto StoneIagw gs=0 TelM="ACQUIRE" PAM- - Adjust CA=-8

Marco reports that shell is not flat, but he is able to set it at 11-12 m/sec local wind.

09:22 pointto StoneIagw gs=0 TelM="ACTIVE" - We can guide, but seeing is too awful to do WFS on this star.

09:28 offset 10 0 (RADEC ABS) - timed out waiting for probe.

09:31 offset -10 0 - worked OK as seen from guider.

Right Side

09:32 Stop PCS, change to Right, Start new PCS

09:33 Register IRTC on Right.

09:34 pointto M5_1332 gs=0 TelM="ACQUIRE" PAM- - Preset succeeded in acquire, but no star. This may be the expected result.

09:38 ptspiral - maybe some pupils went buy

09:40 pointto Arcturus ACQUIRE

09:43 pointto Arcturus TRACK

ptspiral from IA=-280

Find star (or some photons)

IA -280.00000 IE -190.00000 NPAE -12.66810 CA 30.00000 AW -6.65080 AN 22.89390 TF 26.40080 HACA2 -4.10620

09:53 takepic 0.06

Hexapod is still at 0,0,0. OSS was still in simulation mode.

edit lbt.conf

oss.side[1].adsc.use_limits bool true oss.side[1].adsc.simulation bool false

10:02 Stop and restart OSS. Home DX M2 hexapod.

10:08 takepic 0.03 this is now at Andrew's collimation position. A slightly better image than 0.0.0

IA      -280.00000
IE      -195.00000
NPAE     -12.66810
CA        60.00000
AW        -6.65080
AN        22.89390
TF        26.40080
HACA2     -4.10620

10:12 Focussing has improved to a 5 arcsec FWHM core with some spectacular tails by lowering M2 Zglobal from +3.085 to 0.585 mm.

10:14 Stuck in Azimuth, Wayne recovers.

Moving DX M2 RX from 130 to 30 moved IE from -195 to -170 (to maintain the same pointing). Continuing to RX=-70 means IE=-147.

irtc.20100320.102506.fits Z=1.585 irtc.20100320.102556.fits Z=-0.5

M2 X=-0.5 Y=-3.63 Z=0.5 RX -100 RY -100

Adjust more

irtc.20100320.103930.fits Z=1.500 irtc.20100320.104058.fits Z=-0.5

M2 X=-0.5 Y=-3.63 Z above RX=-400 RY=-480
IA      -280.00000
IE       -70.00000
NPAE     -12.66810
CA       -65.00000
AW        -6.65080
AN        22.89390
TF        26.40080
HACA2     -4.10620

takeseq 0.001 100 - irtc.20100320.104225.fits cube 002 Z=-0.5 mm takeseq 0.001 100 - irtc.20100320.104351.fits cube 003 Z=1.5 mm

10:46 pointto M5_1396 TelM="TRACK" PAM- EL=81.6 Focus here is around M2 Zglobal of +0.9. And the aberrations are now comparable to the 3 arcsec seeing.

offset 10 0 RADEC offset 0 0 RADEC offset 10 0 DETXY went along the columns when PCS was expecting rows This is because of the IRTC mounting. offset -10 0 DETXY - strong overshoot, but well aligned with columns (as best you can tell with 3 arcsec seeing)

10:54 stuck in Azimuth

10:56 pointto M5_1396 TelM="TRACK" PAM- - Adjust IE=-74 CA=-66

10:58 pointto M5_1396 gs=0 TelM="GUIDE" - M5 star was too bright for guider - badly saturated

11:00 pointto BS9153 gs=0 TelM="GUIDE" - it is an ugly star, but sextractor should have found it. EL=60

GCS sextractor class is already set to zero, so that isn't the explanation.

11:07 pointto BS9153 ACQUIRE PAM- to focus here on IRTC Zglobal=1.4

11:11 pointto BS9153 gs=0 TelM="GUIDE" -

Changed left to right in right_gcs_sx.sex.

11:18 pointto BS9153 gs=0 TelM="GUIDE" - We are guiding - albeit on a 4 arcsec image. Partly bad seeing and partly aberration from the shell.

11:23 pointto BS9153 gs=8 TelM="GUIDE PA=0 (accidentally)- couldn't move probe for PA=0

11:25 pointto BS9153 gs=8 TelM="GUIDE PA=INDEF=239 - guiding is working off-axis.

takepic 2.0

offset 10 0 RADEC ABS

takepic 2.0 irtc.20100320.113200.fits

offset 0 10 RADEC ABS

takepic 2.0 irtc.20100320.113342.fits

11:40 dropped to none mode

takepic 2.0 empty field irtc.20100320.114103.fits

11:41 pointto BS9153 gs=8 TelM="GUIDE" PA=INDEF=239 - Fails - Rotator not moveable. Apparently Rotator dropped to velocity mode. Cause is not known. See Issue 2628.

11:43 pointto BS9153 gs=8 TelM="GUIDE" PA=INDEF=239 - image looks like a Romulan warbird, but we are guiding again.

11:45 pointto BS9153 gs=10 TelM="GUIDE" PA=INDEF=187.83 - failed not stable (legitimate given the image)

11:49 pointto BS9153 gs=8 TelM="GUIDE" PA=INDEF=239

11:50 pointto BS9153 gs=8 TelM="ACTIVE" PA=INDEF=239

Many of these messages when we have an ACTIVE preset. They are swamping LSS. This is Issue 2619.
Sat Mar 20 11:52:23.594 2010  setting active mode on right side
Sat Mar 20 11:52:24.160 2010  setting active mode on right side
Sat Mar 20 11:52:24.724 2010  setting active mode on right side
Sat Mar 20 11:52:25.290 2010  setting active mode on right side
Sat Mar 20 11:52:25.854 2010  setting active mode on right side
Sat Mar 20 11:52:26.419 2010  setting active mode on right side

12:02 pointto M5_0593 gs=0 TelM="ACQUIRE" - This image is like we had before (not a Romulan warbird). at EL=79.5
IA      -280.00000
IE       -70.00000
NPAE     -12.66810
CA       -62.00000
AW        -6.65080
AN        22.89390
TF        26.40080
HACA2     -4.10620

12:08 pointto M5_0593 gs=2 TelM="ACTIVE" PA=INDEF=274 - Acquisition offset failed.

12:11 pointto M5_0593 gs=2 TelM="TRACK" (decently centered on IRTC)

12:13 pointto M5_0593 gs=2 TelM="ACQUIRE" - This preset hung on the collimating M2 step.

12:17 Stop and Restart OSS. Didn't help - how do I eliminate this stuck preset? Would OSS have timed out?

12:20 Stop IIF, Kill IIF, reregister

12:21 pointto M5_0593 gs=2 TelM="ACQUIRE"

PresetTelescope result status: Error right Preset busy (maybe IIF didn't die)

Killed IIF again and verified that process was gone. So who made this right Preset busy message?

12:25 kill GCSR and restart
Sat Mar 20 12:25:18.498 2010  Preset command busy for right IRTC from IIF
Sat Mar 20 12:26:05.306 2010  Preset command busy for right IRTC from IIF

Reauthorizing clears the preset.

12:28 pointto M5_0593 gs=2 TelM="ACQUIRE" - hung again waiting for M2, We can't understand what it stuck.

12:31 stop PSFR, kill PSFR,

12:33 Collimate of M2 fails to return. See Issue 2620.


12:35 Close for Stuck M2 and impending dawn T=8 degC D=-25.1 degC Sky is photometric.

-- JohnHill - 20 Mar 2010
I Attachment Action Size Date Who Comment
221f9a7.jpgjpg 221f9a7.jpg manage 162 K 23 Mar 2010 - 10:23 JohnHill Actual Romulan warbird for calibration purposes
acquire000009_20100320.jpgjpg acquire000009_20100320.jpg manage 107 K 23 Mar 2010 - 10:21 JohnHill Acquisition image from 11:43 UT that looks like a Romulan warbird
Topic revision: r6 - 23 Mar 2010, JohnHill
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback